Osei is 12 years old and his father is 38years.
(i) In how many years will the father be twice as old as OSei?
(ii)What would be Osei's age then?

Answer by macston(5194)

.
O=Osei's age=12yrs; F=Father's age=38yrs; x=years until father is twice as old
.
F+x=2(O+x) Plug in values.
38yrs+x=2(12yrs+x)
38yrs+x=24yrs+2x Subtract (24yrs+x) from each side.
14yrs=x
ANSWER 1: In 14 years, Father will be twice as old as Osei
ANSWER 2: In 14 years, Osei will be (12yrs+14yrs) 26 years old,
and Father will be (38yrs+14yrs) 52 years old.